# Reward-Aware Proto-Representations in Reinforcement Learning

### v1

@@ 1 Introduction
Learning appropriate representations is a key challenge in reinforcement learning (RL). The successor representation (SR) [7], which represents states as the expected discounted number of visits to successor states, is a particularly promising idea. It has been shown to be an effective distance metric for reward shaping [42, 38], a promising inductive bias for temporally-extended exploration [17, 21], and an effective representation for credit assignment [23, 19], and generalization [16]. Furthermore, recent evidence suggests that the SR is a good computational model for explaining decision-making [24] and neural activity in the brain [33, 32].
The underlying idea behind the SR is to incorporate the temporal aspect of the problem into the representation so that two states are considered similar if they lead to similar future outcomes. Similar ideas permeate other concepts such as proto-value functions [23, 22] and slow feature analysis [31, 41]. In this work, we use the term proto-representations to refer to such temporal representations that implicitly capture the environment dynamics, with the SR the most prominent of them. However, the SR (and others) only captures the transition dynamics of the environment, and does not take rewards into account. In this paper, we demonstrate the benefits of learning a proto-representation that also takes rewards into account.
@@
In this paper, we (1) derive methods for learning the DR using dynamic programming (DP) and temporal-difference (TD) learning [34], and prove the convergence for the DP method, (2) characterize the basis for the vector space of the DR, also describing the settings in which the DR and the SR have the same eigenvectors, (3) define default features, taking a first step towards using the DR with function approximation, and (4) extend the ideas behind the DR and the SR to the maximum entropy RL framework.
@@ 2.1 Reinforcement Learning and the Successor Representation
In this framework, the successor representation [SR; 7] represents states as the expected discounted number of visits to their successor states. Given a policy $\pi$, the SR, $\mathbf{\Psi}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$, is defined as
@@
where $\mathbbm{1}$ is the indicator function. Furthermore, denoting the transition probability matrix induced by $p$ and $\pi$ by $\mathbf{P}^{\pi}$, the SR can be computed in closed form by $\sum_{t=0}^{\infty}\gamma^{t}({\mathbf{P}}^{\pi})^{t}=(\mathbf{I}-\gamma\mathbf{P}^{\pi})^{-1}$. Note that, if we use $\tau_{s\to s^{\prime}}$ to denote a trajectory from $s$ to $s^{\prime}$, we can express the entries of the SR as
${\mathbf{\Psi}}^{\pi}(s,s^{\prime})=\sum_{\tau_{s\to s^{\prime}}}{\mathbf{P}}^{\pi}(\tau_{s\to s^{\prime}})\gamma^{\eta(\tau_{s\to s^{\prime}})},$ (2)
@@
Finally, the SR can be learned using temporal-difference learning [34], for all $j\in\mathcal{S}$:
@@ 2.2 Linearly Solvable Markov Decision Processes
The default representation (DR) is defined in the framework of linearly solvable MDPs [36, 37], a simplification of MDPs in which the optimal value function can be expressed by a linear equation. In this work, building on the work by Piray and Daw [27], we define linearly solvable MDPs as $(\mathcal{S},\mathcal{A},r,p,\pi_{d})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $r:\mathcal{S}\to\mathbb{R}$ is the reward function, $p:\mathcal{S}\times\mathcal{A}\to{\Delta(\mathcal{S})}$ is the transition function, and $\pi_{d}$ is the default policy. It is standard in linearly solvable MDPs to assume no discounting of rewards, i.e., $\gamma=1$. In order to make MDPs linearly solvable, the agent is tasked not only with maximizing the rewards, but also with not deviating too much from the default policy. Deviating from the default policy incurs a cost to the agent, causing the agent to also receive a penalty $\lambda\operatorname{KL}(p^{\pi}\|p^{\pi_{d}}|S_{t})$ at every time step $t$, where $\operatorname{KL}(u\|v)$ denotes Kullback-Leibler (KL) divergence between $u$ and $v$, $p^{\pi}(s^{\prime}|s)$ denotes the probability of transitioning to $s^{\prime}$ from $s$ under policy $\pi$, and $\lambda>0$ determines the relative importance of the deviation cost.
The DR, introduced under this formulation, is a reward-aware proto-representation that encodes the expected rewards for visiting successor states. Let $\mathbf{r}\in\mathbb{R}^{|\mathcal{S}|}$ be the vector of rewards at all states, and $\mathbf{P}^{\pi_{d}}$ be the transition probability matrix induced by $\pi_{d}$ (i.e., $\mathbf{P}^{\pi_{d}}(s,s^{\prime})$ is the probability of transition to $s$ from $s^{\prime}$ under $\pi_{d}$), the DR, denoted by $\mathbf{Z}$, can be computed in closed form by
@@ 3.1 The Default Representation as a Generalization of the Successor Representation
Figure 1: Episodic envs. adapted to incorporate negative rewards. Clockwise: 1) grid task [7], 2) four rooms [35], 3) grid room [38], and 4) grid maze [38]. Start state is in blue. The agent receives $-1$ reward at every time step unless it steps on red tiles ($-20$ reward) or reaches the goal in green (00 reward).
@@
Suppose both the SR and DR are computed with respect to the same policy, i.e., $\pi=\pi_{d}$. When the reward function is constant, i.e., $r(s)=r(s^{\prime})\ \forall s,s^{\prime}\in\mathcal{S}$, the $i$-th eigenvectors of the SR and DR are equivalent, and the $i$-th eigenvalues of the SR ($\lambda_{\text{SR},i}$) and DR ($\lambda_{\text{DR},i}$) are related as follows:
$\lambda_{\text{SR},i}=[\gamma(\lambda_{\text{DR},i}^{-1}-\exp(-r(s)/\lambda)+\gamma^{-1})]^{-1},$ (6)
where $\gamma$ is the discount factor of the SR, $r(s)$ is the state reward, and $\lambda$ is the relative importance of the deviation cost of the DR.
Proof sketch. We write the DR as $[\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]^{-1}$, where $\tilde{r}=\exp(-r(s)/\lambda)$, and perform algebraic manipulations on the expression $\mathbf{Z}\mathbf{e}=\lambda\mathbf{e}$ to obtain the eigenvectors of $[{\mathbf{I}}-\gamma{\mathbf{P}}^{\pi}]^{-1}$. The complete proof is available in Appendix B.2.
This result is particularly interesting because the eigenvectors of the SR have been shown to be useful in settings such as reward shaping and option discovery, which we revisit in Section 6.
@@
${\mathbf{Z}}(s,s^{\prime})=\sum_{\tau_{s\to s^{\prime}}}{\mathbf{P}}^{\pi_{d}}(\tau_{s\to s^{\prime}})\exp(r(\tau_{s\to s^{\prime}})/\lambda),$ (7)
@@ 3.2 State-Action-Dependent Rewards
To extend the DR to state-action pairs, we formulate the cost of deviating from the default policy, $\pi_{d}$, in a way that allows the Bellman optimality equation for the state-action value function to be expressed linearly. We provide the complete derivation in Appendix B.3. Interestingly, the resulting equation looks similar to Equation 4. Using the overbar for vectors and matrices in the state-action setting, e.g., $\bar{{\mathbf{P}}}$ instead of ${\mathbf{P}}$, the DR for state-action-dependent rewards is:
@@
As we show in Section 5.1, the DR under this formulation, combined with default features, unlike prior work, allows direct computation of optimal policies when terminal rewards change, enabling efficient transfer without assuming access to transition dynamics [27, 2].
@@ 4.1 Dynamic Programming
Let $\mathcal{S}_{N}$ denote the set of non-terminal states. Assume $r(s)<0\ \forall s\in\mathcal{S}_{N}$. Let ${\mathbf{R}}=\operatorname{diag}\big{(}\exp(-{\mathbf{r}}/\lambda)\big{)}$, where ${\mathbf{r}}$ is the vector of all state rewards. Let ${\mathbf{Z}}_{0}={\mathbf{R}}^{-1}$. The update rule
@@
Proof sketch. We recursively expand the RHS of Equation 11 and look at its closed-form in the limit, which we show is a Neurmann series. We conclude the proof by leveraging the convergence of the Neumann series and plugging it back into the derived limit. The complete proof is in Appendix B.4.
@@ 5 Extending The DR Beyond Classical Tabular RL
We have been using the tabular case to formalize the ideas we are introducing; now we extend them beyond this setting. Given Equation 13, it is trivial to formalize the parameterization of the DR such that ${\mathbf{Z}}(s,s^{\prime};\bm{\theta})\approx{\mathbf{Z}}(s,s^{\prime})$, for some parameters $\bm{\theta}$. In this section, we go beyond that, defining the concept of default features, which is more semantically meaningful when talking about the DR in terms of features. We also define a proto-representation for the maximum entropy RL framework.
5.1 Default Features
Similar to successor features [SFs; 1], we derive here a decomposition of the value function into features and weights such that these features capture the dynamics of the environment. Different from SFs, default features also consider non-terminal rewards as part of the environment dynamics.
@@ 5.1 Default Features
where ${\mathbf{\Phi}}\in\mathbb{R}^{|S_{T}|\times d}$ is the matrix of terminal state features. We define the product ${\mathbf{Z}}_{NN}{\mathbf{P}}^{\pi_{d}}_{NT}{\mathbf{\Phi}}$ as the DFs matrix, and each row $\bm{\zeta}(s)=({\mathbf{Z}}_{NN}{\mathbf{P}}^{\pi_{d}}_{NT}{\mathbf{\Phi}})_{s,:}$ to be the DFs of a non-terminal state. Then, the optimal value for each non-terminal state can be computed as $\bm{\zeta}(s)^{\top}{\mathbf{w}}$.
@@
where $\alpha$ is the step size, and we define $\bm{\zeta}(s)\coloneqq\bm{\phi}(s)$ for terminal states, $s$. DFs can be thought of as features propagating from the terminal states to all non-terminal states in a manner that respects the reward and transition dynamics. The definition of DFs can be easily extended to the case of state-action-dependent rewards, following our formulation in Section 3.2.
Apart from enabling function approximation, another benefit of learning the DFs over the DR is that it does not require access to the transition dynamics, ${\mathbf{P}}^{\pi_{d}}_{NT}$, to retrieve optimal values. This is especially useful when learning the DFs for state-action-dependent rewards, since we can directly retrieve the optimal policy using the corresponding state-action values (see Equation 10). After learning the DFs in this setting, even without access to environment dynamics, we can efficiently compute the optimal policy under any terminal state reward configuration, enabling transfer learning, as shown in Section 6.
5.2 Proto-Representation in Maximum Entropy RL
Definition 5.1.
We provide the full derivation in Appendix B.5. The MER differs from the DR (Equation 4) in that it uses the adjacency matrix rather than the default policy’s transition probabilities. This connection arises because maximum entropy RL and linearly solvable MDPs are closely related as one can be transformed into the other [8]. As a result, most of our contributions readily extend to the MER. Preliminary analyses showed that the MER also behaves similarly to the DR, so we focus our experiments on the DR. Still, the MER and DR are distinct proto-representations with different formulations, assumptions, and value functions, which can lead to different optimal policies. While we did not observe substantial differences in the settings we considered, exploring them in other regimes remains an open direction.
@@ 6.1 Reward Shaping
Assuming ${\mathbf{e}}$ is the top eigenvector of the SR, prior work [42, 38] has used the shaping reward $\hat{r}_{t}=-\big{(}{\mathbf{e}}(s_{\text{goal}})-{\mathbf{e}}(s_{t+1})\big{)}^{2}$, where ${\mathbf{e}}(s)$ denotes the entry for state $s$. Here, we instead use the top eigenvector ${\mathbf{e}}$ of the DR for potential-based reward shaping [25], defining the shaping reward as $\hat{r}_{t}=\gamma{\mathbf{e}}(s_{t+1})-{\mathbf{e}}(s_{t})$. We compare our approach (DR-pot) with three baselines: 1) potential-based shaping with the SR’s top eigenvector (SR-pot), 2) the distance-based SR method from [38] (SR-pri), and 3) no shaping (ns).
We carry out the experiments in the environments shown in Figure 1. Importantly, the shortest path from the start state to the goal state passes through low-reward regions and is sub-optimal, so the agent needs to learn to avoid low-reward regions in the environment to achieve the optimal return. For the reward shaping approaches, we train a Q-learning [40] agent using a convex combination of the original environment reward, $r_{t}$, and the shaping reward, $\hat{r}_{t}$, resulting in the expression $(1-\beta)r_{t}+\beta\hat{r}_{t}$, where $\beta\in[0,1]$ is a hyperparameter. For the no shaping baseline, we simply train the Q-learning agent using the original environment reward. We use $\gamma=0.99$, $\epsilon=0.05$ for $\epsilon$-greedy exploration, $\lambda=1.3$ for the DR, and perform a grid search over the Q-learning’s step size ($[0.1,0.3,1.0]$) and $\beta$ ($[0.25,0.5,0.75,1.0]$). We run 20 seeds for each hyperparameter setting, and after identifying the best hyperparameters, re-run 50 seeds to avoid maximization bias.

### v2

@@ 1 Introduction
Learning appropriate representations is a key challenge in reinforcement learning (RL). The successor representation (SR) [10], which represents states as the expected discounted number of visits to successor states, is a particularly promising idea. It has been shown to be an effective distance metric for reward shaping [54, 50], a promising inductive bias for temporally-extended exploration [25, 21], and an effective representation for credit assignment [26, 24], generalization [20], and zero-shot RL [47, 48, 6, 44, 1]. Furthermore, recent evidence suggests that the SR is a good computational model for explaining decision-making [28] and neural activity in the brain [41, 40].
The underlying idea behind the SR is to incorporate the temporal aspect of the problem into the representation so that two states are considered similar if they lead to similar future outcomes. Similar ideas permeate other concepts such as proto-value functions [26, 27] and slow feature analysis [39, 53]. In this work, we use the term proto-representations to refer to such temporal representations that implicitly capture the environment dynamics, with the SR the most prominent of them. However, the SR (and others [30, 29, 5, 12, 52]) only captures the transition dynamics of the environment, and does not take rewards into account. In this paper, we demonstrate the benefits of learning a proto-representation that also takes rewards into account.
@@
In this paper, we (1) derive methods for learning the DR using dynamic programming (DP) and temporal-difference (TD) learning [43], and prove the convergence for the DP method, (2) characterize the basis for the vector space of the DR, also describing the settings in which the DR and the SR have the same eigenvectors, and (3) define default features, taking a first step towards using the DR with function approximation.
@@ 2.1 Reinforcement Learning and the Successor Representation
In this framework, the successor representation [SR; 10] represents states as the expected discounted number of visits to their successor states. Given a policy $\pi$, the SR, $\mathbf{\Psi}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$, is defined as
@@
where $\mathbbm{1}$ is the indicator function. Furthermore, denoting the transition probability matrix induced by $p$ and $\pi$ by $\mathbf{P}^{\pi}$, the SR can be computed in closed form by $\sum_{t=0}^{\infty}\gamma^{t}({\mathbf{P}}^{\pi})^{t}=(\mathbf{I}-\gamma\mathbf{P}^{\pi})^{-1}$. Note that in this work, we set the rows corresponding to terminal states to be all zeros for $\mathbf{P}^{\pi}$. Denoting the set of trajectories from $s$ to $s^{\prime}$ by $\mathcal{T}_{s\to s^{\prime}}$, we can express the entries of the SR [5] as
${\mathbf{\Psi}}^{\pi}(s,s^{\prime})=\sum_{{\tau\in\mathcal{T}_{s\to s^{\prime}}}}{\mathbf{P}}^{\pi}(\tau)\gamma^{\eta(\tau)},$ (2)
@@
Finally, the SR can be learned using temporal-difference learning [43, 10], for all $j\in\mathcal{S}$:
@@ 2.2 Linearly Solvable Markov Decision Processes
The default representation (DR) is defined in the framework of linearly solvable MDPs [45, 46], a simplification of MDPs in which the optimal value function can be expressed by a linear equation. In this work, building on the work by Piray and Daw [34], we define linearly solvable MDPs as $(\mathcal{S},\mathcal{A},r,p,\pi_{d})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $r:\mathcal{S}\to\mathbb{R}$ is the reward function, $p:\mathcal{S}\times\mathcal{A}\to{\Delta(\mathcal{S})}$ is the transition function, and $\pi_{d}$ is the default policy that assigns non-zero probabilities to all state-action pairs. It is standard in linearly solvable MDPs to assume no discounting of rewards, i.e., $\gamma=1$. We also make the assumption that $r(s)<0$ for all non-terminal $s$. In order to make MDPs linearly solvable, the agent is tasked not only with maximizing the rewards, but also with not deviating too much from the default policy. Deviating from the default policy incurs a cost to the agent, causing the agent to also receive a penalty $\lambda\operatorname{KL}(p^{\pi}(\cdot|S_{t})\|p^{\pi_{d}}(\cdot|S_{t}))$ at every time step $t$, where $\operatorname{KL}(u\|v)$ denotes Kullback-Leibler (KL) divergence between $u$ and $v$, $p^{\pi}(s^{\prime}|s)$ denotes the probability of transitioning to $s^{\prime}$ from $s$ under policy $\pi$, and $\lambda>0$ determines the relative importance of the deviation cost.
The DR, introduced under this formulation, is a reward-aware proto-representation that encodes the expected rewards for visiting successor states. Let $\mathbf{r}\in\mathbb{R}^{|\mathcal{S}|}$ be the vector of rewards at all states, and $\mathbf{P}^{\pi_{d}}$ be the transition probability matrix induced by $\pi_{d}$ (i.e., $\mathbf{P}^{\pi_{d}}(s,s^{\prime})$ is the probability of transition from $s$ to $s^{\prime}$ under $\pi_{d}$), the DR, denoted by $\mathbf{Z}$, can be computed in closed form by
@@ 3.1 The Default Representation as a Generalization of the Successor Representation
Figure 1: Episodic envs. adapted to incorporate negative rewards. Clockwise: 1) grid task [10], 2) four rooms [42], 3) grid room [50], and 4) grid maze [50]. Start state is in blue. The agent receives 1-1 reward at every time step unless it steps on red tiles (20-20 reward) or reaches the goal in green (0 reward).
@@
Suppose both the SR and DR are computed with respect to the same policy, i.e., $\pi=\pi_{d}$. When the reward function is constant and negative, i.e., $r(s)=r(s^{\prime})<0\ \forall s,s^{\prime}\in\mathcal{S}$, the SR and DR share the same set of eigenvectors. Furthermore, when the SR and DR are symmetric, the $i$-th eigenvectors of the SR and DR are equivalent, and the $i$-th eigenvalues of the SR ($\mu_{\text{SR},i}$) and DR ($\mu_{\text{DR},i}$) are related as follows:
$\mu_{\text{SR},i}=\bigg[\gamma\Big(\mu_{\text{DR},i}^{-1}-\exp\big(-r(s)/\lambda\big)+\gamma^{-1}\Big)\bigg]^{-1},$ (6)
where $\gamma\in(0,1)$ is the discount factor of the SR, $r(s)$ is the state reward, and $\lambda$ is the relative importance of the deviation cost of the DR.
Proof sketch. We write the DR as $\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]^{-1}$, where $\tilde{r}=\exp(-r(s)/\lambda)$, and perform algebraic manipulations on the expression $\mathbf{Z}\mathbf{e}=\lambda\mathbf{e}$ to obtain the eigenvectors of ${\mathbf{I}}-\gamma{\mathbf{P}}^{\pi}]^{-1}$. Then, we show that the eigenvalues of the SR and DR are related by a function that is monotonically increasing over the range of DR eigenvalues. The complete proof is available in Appendix B.2.
This result is particularly interesting because the eigenvectors of the SR have been shown to be useful in settings such as reward shaping and option discovery. The top eigenvector of the SR captures temporal distance between states, and such distance has been used for reward shaping as it leads to better performance than distances that are based on the coordinates of states  [50, 54]. In option discovery, it has been shown that eigenoptions, which are obtained by training a policy to maximize a reward signal induced by the eigenvectors of the SR, allow the agent to reach less explored areas of the state space, facilitating exploration [21]. We will revisit these applications in Section 6.
@@
${\mathbf{Z}}(s,s^{\prime})=\sum_{{\tau\in\mathcal{T}_{s\to s^{\prime}}}}{\mathbf{P}}^{\pi_{d}}(\tau)\exp(r(\tau)/\lambda),$ (7)
@@ 3.2 State-Action-Dependent Rewards
To extend the DR to state-action pairs, we need to extend linearly solvable MDPs to state-action pairs, which was first presented in the work by Ringstrom et al. [37]. However, the DR was not discussed in this work, so we provide our derivation in Appendix B.3. Interestingly, the resulting equation looks similar to Equation 4. Using the overbar for vectors and matrices in the state-action setting, e.g., $\bar{{\mathbf{P}}}$ instead of $\mathbf{P}$}, the DR for state-action-dependent rewards is:
@@
As we show in Section 5, the DR under this formulation, combined with default features, unlike prior work, allows direct computation of optimal policies when terminal rewards change, enabling efficient transfer without assuming access to transition dynamics [34, 3].
@@ 4.1 Dynamic Programming
[…]
@@
Proof sketch. We recursively expand the RHS of Equation 11 and look at its closed-form in the limit, which we show is a Neumann series. We conclude the proof by leveraging the convergence of the Neumann series and plugging it back into the derived limit. The complete proof is in Appendix B.4.
@@ 5 Extending The DR Beyond Classical Tabular RL
We have been using the tabular case to formalize the ideas we are introducing; now we extend them beyond this setting. Given Equation 13, it is trivial to formalize the parameterization of the DR such that ${\mathbf{Z}}(s,s^{\prime};\bm{\theta})\approx{\mathbf{Z}}(s,s^{\prime})$, for some parameters $\bm{\theta}$. In this section, we go beyond that, defining the concept of default features, which is more semantically meaningful when talking about the DR in terms of features. We also define a proto-representation for the maximum entropy RL framework. However, since this paper focuses on the DR, we defer the discussion of this proto-representation to Appendix B.5.
Similar to successor features [SFs; 2], we derive here a decomposition of the value function into features and weights such that these features capture the dynamics of the environment. Different from SFs, default features also consider non-terminal rewards as part of the environment dynamics.
@@
where ${\mathbf{\Phi}}\in\mathbb{R}^{|S_{T}|\times d}$ is the matrix of terminal state features. We define the product ${\mathbf{Z}}_{NN}{\mathbf{P}}^{\pi_{d}}_{NT}{\mathbf{\Phi}}$ as the DFs matrix, and each row $\bm{\zeta}(s)=({\mathbf{Z}}_{NN}{\mathbf{P}}^{\pi_{d}}_{NT}{\mathbf{\Phi}})_{s,:}$ to be the DFs of a non-terminal state. Then, the optimal value for each non-terminal state can be computed as $\bm{\zeta}(s)^{\top}{\mathbf{w}}$. Note that such a form of optimal value computation is only possible when only the rewards at terminal states change. The reason is that, when only the terminal state rewards change, $\zeta(s)$ stays the same for all non-terminal $s$. However, when the rewards for non-terminal states change, $\zeta(s)$ changes since $\zeta(s)=\mathbf{Z}_{NN}\mathbf{P}^{\pi_{d}}_{NT}\mathbf{\Phi}$ and $\mathbf{Z}_{NN}$ depends on non-terminal state rewards.
@@
where $\alpha$ is the step size, and we define $\bm{\zeta}(s)\coloneqq\bm{\phi}(s)$ for terminal states, $s$. DFs can be thought of as features propagating from the terminal states to all non-terminal states in a manner that respects the reward and transition dynamics. The definition of DFs can be easily extended to the case of state-action-dependent rewards, following the formulation in Section 3.2.
Apart from enabling function approximation, another benefit of learning the DFs over the DR is that it does not require access to the transition dynamics, ${\mathbf{P}}^{\pi_{d}}_{NT}$, to retrieve optimal values. This is especially useful when learning the DFs for state-action-dependent rewards, since we can directly retrieve the optimal policy using the corresponding state-action values (see Equation 10). After learning the DFs in this setting, even without access to environment dynamics, we can efficiently compute the optimal policy under any terminal state reward configuration, enabling transfer learning, as shown in Section 6. Note that while DFs enable directly computing the optimal policy, they are limited to the scenario where only the terminal state rewards change. On the other hand, while SFs can only compute a policy as good as the ones used to compute the SFs, they can be applied when any state rewards change, allowing more flexibility than DFs.
@@ 6.1 Reward Shaping
Assuming $\mathbf{e}$} is the top eigenvector of the SR, prior work [54, 50] has used the shaping reward $\hat{r}_{t}=-\big({\mathbf{e}}(s_{\text{goal}})-{\mathbf{e}}(s_{t+1})\big)^{2}$, where ${\mathbf{e}}(s)$ denotes the entry for state $s$. Here, we instead use the top eigenvector $\mathbf{e}$} of the DR for potential-based reward shaping [31], defining the shaping reward as $\hat{r}_{t}=\gamma{\mathbf{e}}(s_{t+1})-{\mathbf{e}}(s_{t})$. We compare our approach (DR-pot) with three baselines: 1) potential-based shaping with the SR’s top eigenvector (SR-pot), 2) the distance-based SR method from [50] (SR-pri), and 3) no shaping (ns). Both the SR and DR are computed with respect to the uniform random policy.
We carry out the experiments in the environments shown in Figure 1. Importantly, the shortest path from the start state to the goal state passes through low-reward regions and is sub-optimal, so the agent needs to learn to avoid low-reward regions in the environment to achieve the optimal return. For the reward shaping approaches, we train a Q-learning [51] agent using a convex combination of the original environment reward, $r_{t}$, and the shaping reward, $\hat{r}_{t}$, resulting in the expression $(1-\beta)r_{t}+\beta\hat{r}_{t}$, where $\beta\in[0,1]$ is a hyperparameter. Note that we assume access to the eigenvectors of the SR and DR prior to training the agent with the potential-based reward. Future work can explore learning the eigenvectors and the policy simultaneously. For the no shaping baseline, we simply train the Q-learning agent using the original environment reward. We use $\gamma=0.99$, $\epsilon=0.05$ for $\epsilon$-greedy exploration, $\lambda=1.3$ for the DR, and perform a grid search over the Q-learning’s step size ($[0.1,0.3,1.0]$) and $\beta$ ($[0.25,0.5,0.75,1.0]$). We run 20 seeds for each hyperparameter setting, and after identifying the best hyperparameters, re-run 50 seeds to avoid maximization bias.